OpenStudy (anonymous):

The gradient at any point (x,y) on a curve is sqrt {1+2x}. The curve passes through the point (4,11). Find (i) the equation of the curve. (ii) the point at which the curve intersects the y-axis.

OpenStudy (anonymous):

Take the anti-derivative of the gradient to find out the equation of the curve. Once the equation has been figured out, put x as zero and get the answer to the second part.

OpenStudy (campbell_st):

(i) requires you to find the primitive or indefinite integral \[f'(x) = ( 1 + 2x)^{1/2}\] then \[F(x) = 1/3(1 + 2x)^{3/2} + c\] substitute x = 4 and F(x) = 11 to evaluate c this will also be the y-intercept

OpenStudy (anonymous):

Won't it be times two, since it's (1 over 2/3)(1+2x)^3/2

OpenStudy (anonymous):

dY/dX=sqrt(1+2x)..integerate both sides Y=1/3(1+2X)^3/2+c.....put the points on this curve and find c constant.. (a)Y=1/3(1+2x)+c...put the value of c. (b) put X=0 in the curve equation and get that point...(0,y)..find Y .

OpenStudy (anonymous):

How do you integrate?

OpenStudy (anonymous):

@order the coefficient of X is 2 the divide by 2 here.

OpenStudy (anonymous):

Why is the final answer not \[{(1+2x)^{3/2} \over {3\over2}}?\]

OpenStudy (anonymous):

@order do you have any problem..

OpenStudy (anonymous):

Yes... How do you integrate it? I did it like this: \[\int\limits 2 + \sqrt 2x^{1/2}\]\[y= x +{ \sqrt2x^{3/2} \over {3\over2}}\]\[y=x+ {2\sqrt2x^{3/2} \over 3}\]

OpenStudy (anonymous):

\[\int\limits 1~~~ sorry\]

OpenStudy (anonymous):

Would that be correct?

OpenStudy (anonymous):

@Taufique

OpenStudy (anonymous):

@FoolForMath

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

here think that you have integrate it with respect to X .not 2X .if you integrate it with respect to 2X then you must divide by 2 here.

OpenStudy (anonymous):

I expanded the bracket though.. Would that be wrong?

OpenStudy (anonymous):

no, it is not wrong.There is many way to integrate any function with respect to variable

OpenStudy (anonymous):

where is your confusion??

OpenStudy (anonymous):

Is my integration correct, or wrong?

OpenStudy (anonymous):

there is mathematical processing error ..so i am unable to read it.

OpenStudy (anonymous):

x + {2{sqrt 2}x^3/2 over 3}

OpenStudy (anonymous):

When you split the square root, it's totally WRONG!

OpenStudy (anonymous):

this is incorect..

OpenStudy (anonymous):

Simple put sqrt ( 4 + 4) not equal to squrt ( 4 ) + sqrt ( 4)

OpenStudy (anonymous):

you cant expand it in this way ..this is wrong way..

OpenStudy (anonymous):

If you expand the brackets, isn't it int sqrt 1 + sqrt 2 sqrtx dx?

OpenStudy (anonymous):

Ok... So, that doesn't work...

OpenStudy (anonymous):

Technically, exponential doesn't work!

OpenStudy (anonymous):

OK. So, can you show me once again, how to integrate it?

OpenStudy (anonymous):

All you do is applying traditional method U substitution!

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

u = √ ( 1 + 2x) -> u² = 1 + 2x => 2udu = 2dx Plug them in!

OpenStudy (kropot72):

Hello order, Your method of integration is not valid. One method to integrate the function is as follows: z = 1+2x dz = 2dx \[\int\limits z ^{1/2}.dz/2 = 1/2\int\limits z ^{1/2}.dz = 1/3(1+2x)^{3/2} + C\]

OpenStudy (anonymous):

As long as you obtain the result (1/3) u^(3/2) + C. That's the correct result!

OpenStudy (anonymous):

hmmm. ok

OpenStudy (anonymous):

you can do direct method: dy=1/2*sqrt(1+2x)d(2x).....2X part is variable integrate both sides y=1/2*((1+2x)^3/2)/(3/2))+c hence y=1/3(1+2x)^3/2+c

OpenStudy (anonymous):

Thanks, I understand now :)

OpenStudy (anonymous):

What's with so much confusion. :O \[\int\limits_{}^{} (1+2x)^{0.5}dx = (1+2x)^{0.5+1}\div0.5+1 + constant\]\[(1+2x)^{1.5} \div 1.5 + constant\] Now use the information given to find out the constant. \[11 = (1+2(4)) \div 1.5 + constant\] The constant comes out to be -7. Therefore, the equation of the curve will be \[y = ((1+2x)^{1.5} \div 1.5) - 7\] To find out the point on the y-axis, put x as zero. The answer comes out to be 3.4641.

OpenStudy (kropot72):

Hello Shahan03, Sorry but your integration is invalid. The basic method for integration cannot be directly used on the particular function to be integrated.

OpenStudy (anonymous):

THE VALUE OF C MUST BE 2..

OpenStudy (anonymous):

kropot72: Why is that so? I have an exam in few days and so your reply will be very helpful.

OpenStudy (anonymous):

Thanks a lot for pointing that out. I was indeed wrong. Apologies for the inconvenience. :)

OpenStudy (kropot72):

To check whether or not your method of integration is valid, differentiate your result with respect to the variable. If your method is correct you will obtain the original function.